OpenStudy (anonymous):

If the equation describing the motion of a rocket is x = 10 + 5t^2, write an expression for the instantaneous speed of the rocket. What is the instantaneous speed at 5 s. What is the acceleration? WHere do I begin? Do I need to take the integral anywhere?

OpenStudy (anonymous):

just derivate equation wrt x to get instantaneous speed a).then put value of t=5 you will have answer b). again derivate wrt x u will have acceleration

OpenStudy (anonymous):

Yes. Velocity = dx/dt acceleration = d^2 x/dt^2

OpenStudy (anonymous):

v(t)=dx(t)/dt and a(t)=dv(t)/dt where x(t)=10+5t^2

OpenStudy (anonymous):

okay...thanks!!! mwah!! :)

OpenStudy (anonymous):

well thank you :)